# Adversarial Robustness of Deep Learning Models for Inland Water Body Segmentation from SAR Images

### v1

@@ 4.4 Observed Anomalies in the Generated Ground Truth
We observed two different types of anomalies in the generated ground truth. Here, we report the findings of the model performance in the present of the anomalies, which may be considered as inadvertent label noise. These anomalies are different from the controlled morphological label noise. Thus, the previous references of “uncorrupted masks” in this paper meant that the masks did not have the controlled label noise, but still contained the inadvertent label noise.
The first anomaly is from the choice of the threshold for the NDWI index when generating the mask for the images. This threshold has been manually selected. At times, this threshold need not be the most optimal one for the image, as shown in Figure 14. As observed, the ground truth is not fully accurate, and contains regions which are not water bodies, but are misclassified as water. However, we observe that the training is robust even in this case, as the mask predicted by the baseline U-Net model in the absence of label noise is qualitatively accurate when compared to the actual image. The same is also observed for all the other models. However, this anomaly can be avoided by adaptively selecting the optimal threshold for the NDWI index [47], while creating the ground truth, i.e., the uncorrupted masks.
Figure 15: Images with accurate masks (green) and anomalous masks (red), where the latter is due to incomplete data in Sentinel-2 images for ground truth generation. Observing these masks in-place in the Amazon River Basin helped identify the cause for this anomaly.
@@
Overall, these results show that all the models trained with uncorrupted masks are fairly robust to these anomalies in the dataset generation, highlighting their suitability for the task of inland water body segmentation. However, the only limitation of these models, as alluded earlier, is in the accurate delineation of small tributaries.
We chose the U-Net model for adversarial experimentation primarily due to its ability to achieve favorable performance even with relatively smaller training sets, and their ability to run faster than other models. Also, U-Net is known to be more stable to errors than other architectures [28]. Future scope of work can include performing similar experiments on other segmentation architectures, and to compare stability and performance. One line of work can be to observe the effects of only erosion and only dilation on the datasets, to quantify and compare the extent of degradation in model performance by the two methods. Another interesting direction of work would be to include other, more traditional modes of corruption, including rotation and translation, and observing their effects.

### v2

@@ 4.4 Observed Anomalies in the Generated Ground Truth
We observed two different types of anomalies in the generated ground truth. Here, we report the findings of the model performance in the presence of the anomalies, which may be considered as inadvertent label noise. These anomalies are different from the controlled morphological label noise. Thus, the previous references of “uncorrupted masks” in this paper meant that the masks did not have the controlled label noise, but still contained the inadvertent label noise.
The first anomaly is from the choice of the threshold for the NDWI index when generating the mask for the images. This threshold has been manually selected. At times, this threshold need not be the most optimal one for the image, as shown in Figure 14. As observed, the ground truth is not fully accurate and contains regions that are not water bodies but are misclassified as water. However, we observe that the training is robust even in this case, as the mask predicted by the baseline U-Net model in the absence of label noise is qualitatively accurate when compared to the actual image. The same is also observed for all the other models. However, this anomaly can be avoided by adaptively selecting the optimal threshold for the NDWI index [48], while creating the ground truth, i.e., the uncorrupted masks.
Figure 15: Images with accurate masks (green) and anomalous masks (red), where the latter is due to incomplete data in Sentinel-2 images for ground truth generation. Observing these masks superimposed on the Amazon River Basin helped identify the cause for this anomaly.
@@
Overall, these results show that all the models trained with uncorrupted masks are fairly robust to these anomalies in the dataset generation, highlighting their suitability for the task of inland water body segmentation. However, the only limitation of these models, as alluded to earlier, is in the accurate delineation of small tributaries.
We chose the U-Net model for adversarial experimentation primarily due to its ability to achieve favorable performance even with relatively smaller training sets, and its ability to run faster than other models. Also, U-Net is known to be more stable to errors than other architectures [29]. Future scope of work can include performing similar experiments on other segmentation architectures and comparing stability and performance. One line of work can be to observe the effects of only erosion and only dilation on the datasets, to quantify and compare the extent of degradation in model performance by the two methods. Another interesting direction of work would be to include other, more traditional modes of corruption, including rotation and translation, and observing their effects.